Calculate the volume of acid needed to prepare 500ml of 2.0 molar concentration of acid,if the stock bottle of the acid reads 1.25 specific gravity, 56% by mass and molar mass of acid 56.


1
Expert's answer
2021-10-22T03:58:48-0400

To calculate the molarity of a solution that is 56% by mass, begin by calculating the total mass of one L of the solution. Given that the specific gravity of the solution is 1.25, we know that one L of the solution weighs 1250g. Multiplying the total weight by the percentage we get the number of grams of acid per liter.

"56.0 \% \times 1250 \;g/L = 700 \;g/L"

Then divide by the molecular weight to obtain the molarity.

"Molarity = 700 \;g/L \times \frac{1\;mol}{56 \;g} = 12.5 \;M"

500 ml of 2.0 M solution will contain 1 mol of acid.

Proportion:

12.5 mol – 1000 mL

1 mol – x

"x = \frac{1 \times 1000}{12.5} = 80 \;mL"

Answer: 80 mL